MILLIMETER-WAVE SPECTROSCOPY OF THE IRON CARBONYL RADICAL(FeCO) IN THE ν_{2} BEND-ING VIBRATIONAL STATE

\underline{KEIICHI TANAKA}, MITSUHIRO NAKAMURA, MITSUAKI SHIRASAKA and TAKEHIKO TANAKA, Department of Chemistry, Faculty of Science, Kyushu University 33, Hakozaki, Higashiku, Fukuoka 812-8581, Japan, and Institute for Molecular Science, Myodaiji, Okazaki 444-8585, Japan.

The rotational spectrum of iron carbonyl radical FeCO in the ν_{2} bending vibrational state of the ground X^{3}Σ^{−} electronic state was observed in the millimeter-wave region. The radical was produced by a dc discharge of iron pentacarbonyl Fe(CO)_{5}. Seven rotational lines, split into sextet by the electron spin-spin interaction and Λ-type dobling, were identified in the frequency region of 154-254 GHz. Molecular constants derived, the rotational and centrifugal distortion constants, the spin-spin coupling constant λ_{0} = 679.37(291)GHz and the spin-rotation coupling constant γ_{0} = −1180.0(19) MHz, are reasonably similar to those of the ground state ^{a}. Although the electronic state is ^{3}Σ^{−}, a large spin-orbit coupling constant A = 14.0746(57) GHz was derived, indicating the vibronic coupling with the ^{3}Π electronic state located about 6500 cm^{−1} above the ground state. From the Λ-type doubling, large interaction constants, o = −18.2899(55) GHz, p = −355.36(64) MHz, and q = 9.5167(69) MHz, were determind, where the figures in parentheses are a standard error to be attached to the last digits.

aK. Tanaka, M. Shirasaka and T. Tanaka, J. Chem. Phys. \underline{106}, 6820 (1997)